# New Results on the Polyak Stepsize: Tight Convergence Analysis and Universal Function Classes

### v2

@@ Abstract
In this paper, we revisit a classical adaptive stepsize strategy for gradient descent: the Polyak stepsize (PolyakGD), originally proposed in polyak1969minimization. We study the convergence behavior of PolyakGD from two perspectives: tight worst-case analysis and universality across function classes. As our first main result, we establish the tightness of the known convergence rates of PolyakGD by explicitly constructing worst-case functions. In particular, we show that the $\mathcal{O}((1-\frac{1}{\kappa})^{K})$ rate for smooth strongly convex functions and the $\mathcal{O}(1/K)$ rate for smooth convex functions are both tight. Moreover, we theoretically show that PolyakGD automatically exploits floating-point errors to escape the worst-case behavior. Our second main result provides new convergence guarantees for PolyakGD under both Hölder smoothness and Hölder growth conditions. These findings show that the Polyak stepsize is universal, automatically adapting to various function classes without requiring prior knowledge of problem parameters.
@@ 1 Introduction
The Polyak stepsize is a stepsize schedule for gradient descent in convex optimization, originally proposed by Boris T. Polyak (polyak1969minimization; polyak1987introduction). It was initially designed for the subgradient method in nonsmooth optimization, and has recently been extended to more general settings (hazan2019revisiting; loizou2021stochastic; wang2023generalized; deng2024uniformly; devanathan2024polyak). The Polyak stepsize requires knowledge of the optimal function value $f^{\star}$ and is defined as
@@
PolyakGD automatically adjusts the stepsize according to the position of the current iterate $x^{k}$ on the optimization landscape. The resulting stepsizes are typically much larger than standard constant or diminishing stepsizes, making PolyakGD more aggressive and often yielding superior practical performance (zamani2024exact). Consequently, the Polyak stepsize is widely used in settings where the optimal function value $f^{\star}$ is known, such as convex feasibility problems and over-parameterized machine learning models (boyd2013subgradient; loizou2021stochastic). Even when $f^{\star}$ is not known a priori, PolyakGD remains highly effective, provided that a reasonable estimate of the optimal value is available (polyak1969minimization; hazan2019revisiting; zheng2024adaptive).
@@
From a theoretical perspective, the convergence of the Polyak stepsize has been extensively studied in nonsmooth convex optimization. polyak1987introduction proved that the subgradient method with the Polyak stepsize satisfies $\lim_{k\to\infty}\sqrt{k}\,(f(x^{k})-f^{\star})=0$. Moreover, the convergence rate of the best-iterate [^1] By best-iterate, we refer to the quantity $\min_{1\leq k\leq K}(f(x^{k})-f^{\star})$; by last-iterate, we refer to $f(x^{K})-f^{\star}$. is $\mathcal{O}(1/\sqrt{K})$ (see, e.g., boyd2013subgradient). Beyond best-iterate guarantees, zamani2024exact analyzed the last-iterate behavior of the Polyak stepsize, establishing an $\mathcal{O}(1/\sqrt[4]{K})$ convergence rate and providing a matching worst-case function. Their analysis uses the performance estimation problem (PEP) methodology (drori2014performance; taylor2017smooth), building on earlier work (zamani2025exact)222zamani2024exact; zamani2025exact employ PEP to obtain numerical dual multipliers that guide the derivation of analytical expressions for their proofs.. Additionally, criscitiello2023curvature provided a worst-case function for subgradient method with the Polyak stepsize on Riemannian manifolds. More recently, orabona2025new showed that the Polyak stepsize can be interpreted as applying the subgradient method to a specific surrogate function, providing an alternative justification for the need to know $f^{\star}$. In contrast to the well-studied nonsmooth setting, the smooth setting remains far less explored. The primary reference is the note by hazan2019revisiting, who established convergence guarantees for PolyakGD on both smooth convex and smooth strongly convex objectives. In the smooth convex setting, they proved an $\mathcal{O}(1/K)$ convergence rate for the best-iterate, matching the rate of constant-stepsize gradient descent.
Although PolyakGD often exhibits superior practical performance, the existing $\mathcal{O}(1/K)$ convergence rate matches that of gradient descent with a constant stepsize. This agreement was unsurprising prior to recent PEP-based convergence analysis (grimmer2024provably; altschuler2025acceleration). For example, grimmer2024provably showed that gradient descent with frequent long steps achieves provably faster convergence. Unfortunately, the PEP methodology does not readily extend to adaptive stepsizes for general convex functions (goujaud2023fundamental; goujaud2024pepit), applying current computer-assisted analysis techniques remains challenging. To our knowledge, the only two existing papers studying the tight convergence of the Polyak stepsize focus on the strongly convex case and only deal with variants of the original Polyak stepsize (barre2020complexity; huang2024analytic). In the strongly convex setting, PEP remains applicable, as one can analyze the contraction ratio between consecutive iterations. This line of work naturally raises the following question:
@@
In this paper, we show that the known convergence rates for PolyakGD are indeed tight by constructing worst-case functions. On the other hand, we demonstrate that under floating-point arithmetic, PolyakGD can automatically leverage numerical error to escape this worst-case behavior. This interesting phenomenon explains the superior empirical performance of PolyakGD in practice. Theoretically, the strong performance of PolyakGD is attributed to its inherent adaptivity, a crucial property for modern optimization methods (malitsky2020adaptive; gao2024gradient; li2025simple; he2025history). hazan2019revisiting showed that the Polyak stepsize is simultaneously adaptive to both smooth and nonsmooth regimes, encompassing general convex as well as strongly convex objectives. mishkin2024directional further established that it adapts to directional smoothness. More recently, orabona2025new demonstrated that the Polyak stepsize also adapts to the curvature of a surrogate function induced by the original objective. Inspired by these results, we further advance this line of research and provide a comprehensive answer to the following question:
@@
Contributions.
This work answers the above two questions. To address the first question, we construct a simple two-dimensional quadratic function. Our proof strategy relies on choosing a specific initial point such that the Polyak stepsize reduces to a constant stepsize along the trajectory in the strongly convex case. Subsequently, we derive the worst-case functions for the general convex and Hölder smooth cases by adapting this quadratic function. Based on this construction, we establish the tightness of the linear convergence rate for strongly convex functions, the $\mathcal{O}(1/K)$ rate for general convex functions, and the $\mathcal{O}(1/K^{(\nu+1)/2})$ rate for $\nu$-Hölder smooth objectives. We then show that under floating-point arithmetic, $\gamma$-PolyakGD escapes this worst-case behavior.
Our second result demonstrates that the Polyak stepsize is universal across various function classes. We establish that PolyakGD simultaneously adapts to both Hölder smoothness and Hölder growth conditions (Theorem 4.1). In particular, when only the Hölder growth condition holds, $\gamma$-PolyakGD achieves the optimal convergence rate; when the function is Hölder smooth, it matches the rate of the universal gradient method of nesterov2015universal. Furthermore, we show that the Polyak stepsize adapts to the recently proposed intrinsic global curvature bound (nesterov2025universal), and that our universal analysis extends to the stochastic setting under the interpolation condition. A summary of these results is provided in Table 1.
@@
hazan2019revisiting Theorem 3.2
@@
hazan2019revisiting Theorem 3.1
@@
Theorem 4.1 nemirovskii1985optimal
@@ 2 Preliminaries
where the objective function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is convex and differentiable [^3] Our analysis can be extended to the projected (sub)gradient method for the constrained setting $\min_{x\in\mathcal{X}}f(x)$ with a closed convex constraint set $\mathcal{X}$ and $f$ is nonsmooth.. We assume that the optimal solution set $\mathcal{X}^{\star}$ is nonempty and the optimal function value $f^{\star}=f(x^{\star})$ is finite. Besides, we assume that $f^{\star}$ is known. Addressing the scenario where $f^{\star}$ is unknown is beyond the scope of this paper, and the techniques developed in polyak1969minimization; hazan2019revisiting can be applied to estimate $f^{\star}$ at the cost of an additional $\log(1/\varepsilon)$ factor in the overall complexity.
@@
with proofs provided in orabona2023normalized. For the lower curvature bound, we adopt the following growth condition:
@@
Consequently, several subtle points deserve attention. First, we require that $\mathcal{X}^{\star}\subseteq\mathcal{K}$, otherwise the Hölder growth condition does not seem intuitive. As a result, we exclude the case $r<\nu+1$, as choosing $x$ sufficiently close to $\mathcal{X}^{\star}$ would lead to a contradiction with the inequality (6). Moreover, we require that $\sup_{x\in\mathcal{K}}\operatorname{dist}(x,\mathcal{X}^{\star})<\infty$. Otherwise, the case $r>\nu+1$ would also yield a contradiction by taking $x$ such that $\operatorname{dist}(x,\mathcal{X}^{\star})\rightarrow\infty$ (see Section 1.2 of roulet2017sharpness for a similar argument). Fortunately, as discussed at the beginning of Section 4, $\gamma$-PolyakGD (Algorithm 1) naturally ensures the existence of such a proper set $\mathcal{K}$ due to the Fejér monotonicity.
@@ 3 Tight Convergence of $\gamma$-PolyakGD
Consequently, the main task is to construct a worst-case function to answer the above question. For instance, drori2014performance; teboulle2023elementary utilized the Huber loss to establish the tightness of the $\mathcal{O}(1/K)$ rate for constant stepsize gradient descent. The intuition is that the small gradient magnitude of the Huber loss forces a constant stepsize gradient descent to take conservatively small steps, thereby justifying the tightness. However, this strategy fails due to the adaptivity of $\gamma$-PolyakGD, as demonstrated below.
@@ 3.1 Worst-case quadratic functions for $\gamma$-PolyakGD
Our reduction argument is also applicable to the BB stepsize on quadratic functions due to its scale invariance. By leveraging the tight convergence result and the worst-case quadratic function from li2021faster, a similar $\Omega(1/K)$ result can be established for the BB stepsize.
@@
Setting $\nu\rightarrow 0$, the above result reduces to the lower bound $\Omega(1/\sqrt{K})$. We remark that this result does not contradict the $\Omega(1/\sqrt[4]{K})$ bound established in zamani2024exact. Our tightness result is established for the best iterate, while theirs applies to the last iterate, which is generally a more challenging criterion. Consequently, the worst-case function constructed in zamani2024exact is more complicated. This distinction reveals the significant gap between the best iterate and the last iterate under bounded gradient condition, highlighting the necessity of analyzing both cases.
@@
Establishing the tight gradient norm convergence of constant stepsize gradient descent involves subtle distinctions. Setting $\nu=1$, the above result reduces to the lower bound $\Omega(1/\sqrt{K})$. However, nesterov2018lectures; diakonikolas2021potential; teboulle2023elementary established that small constant stepsizes achieve an upper bound of $\mathcal{O}(1/K)$, a result shown to be tight by kim2018generalizing via PEP. Using a slightly large constant stepsize slows the convergence to $\mathcal{O}(1/\sqrt{K})$, as shown in Proposition 4.4 in rotaru2022tight and Theorem 2.1 in rotaru2024exact, a bound that is also tight according to PEP. Therefore, our result suggests that the Polyak stepsize behaves like a “large constant stepsize” in some sense.
@@ 3.2 Polyak stepsize escapes the worst case under floating-point arithmetic
Our tightness convergence results show that PolyakGD reduces to gradient descent with a constant stepsize on the worst-case quadratic function. While this conclusion holds under exact arithmetic, it may not persist in practical implementations due to floating-point errors. Indeed, when implementing the Polyak stepsize on the constructed worst-case quadratic function, the numerical behavior differs from the theoretical results. Figure 2 plots the stepsize sequence $\{\alpha_{k}\}$ and the corresponding function value gaps $\{q(x^{k})-q^{\star}\}$ generated by $1$-PolyakGD. As predicted by our theory, $1$-PolyakGD initially behaves like constant stepsize gradient descent. Yet after a few iterations, it departs from the theoretical trajectory, producing larger stepsizes and converging significantly faster. This suggests that under inexact arithmetic, $\gamma$-PolyakGD (with $\gamma\in(0,2)$) can automatically exploit floating-point errors to escape the worst-case behavior, which may explain the superior practical performance of the Polyak stepsize. This section formally quantifies this intuition. The techniques developed in this section have also been applied to analyze the hypergradient descent heuristic recently (chu2025gradient).
@@
and $s_{1}$ (or equivalently $s_{2}$) is a period-2 orbit of $\mathcal{F}_{\gamma}$ (alligood1997chaos). To understand the local behavior of $\gamma$-PolyakGD in the presence of floating-point errors, it is natural to examine the stability of this nonlinear map around its orbit. This reduces to analyzing the spectrum of the Jacobian of $\mathcal{F}_{\gamma}$. The proof of Proposition 3.2 is assisted by Mathematica and the source code is available at this link.
@@
Proposition 3.2 shows that when $\gamma=1$ and $\kappa>6$, the spectral radius of the product Jacobian is strictly greater than $1$. Hence, the nonlinear dynamical system (13) is unstable around the trajectory induced by the worst-case function. Consequently, once floating-point errors enter the iterations, $\gamma$-PolyakGD deviates from this trajectory and exhibits accelerated convergence. Similar observation holds for all $\gamma\in(0,2)$. In contrast, when $\gamma=2$, the spectral radius is exactly $1$, implying that $2$-PolyakGD may become trapped on the worst-case trajectory. It is consistent with the zig-zag behavior observed in huang2024analytic. In summary, the worst-case trajectory is only realized under exact arithmetic, and PolyakGD in practice is largely immune to it.

### v3

@@ Abstract
In this paper, we revisit a classical adaptive stepsize strategy for gradient descent: the Polyak stepsize (PolyakGD), originally proposed in Polyak (1969). We study the convergence behavior of PolyakGD from two perspectives: tight worst-case analysis and universality across function classes. As our first main result, we establish the tightness of the known convergence rates of PolyakGD by explicitly constructing worst-case functions. In particular, we show that the $\mathcal{O}((1-\frac{1}{\kappa})^{K})$ rate for smooth strongly convex functions and the $\mathcal{O}(1/K)$ rate for smooth convex functions are both tight. Moreover, we theoretically show that PolyakGD automatically exploits floating-point errors to escape the worst-case behavior. Our second main result provides new convergence guarantees for PolyakGD under both Hölder smoothness and Hölder growth conditions. These findings show that the Polyak stepsize is universal, automatically adapting to various function classes without requiring prior knowledge of problem parameters.
@@ 1 Introduction
The Polyak stepsize is a stepsize schedule for (sub)gradient descent in convex optimization, originally proposed by Boris T. Polyak (Polyak, 1969, 1987). It was initially designed for the subgradient method in nonsmooth optimization, and has recently been extended to more general settings (Hazan and Kakade, 2019; Loizou et al., 2021; Wang et al., 2023; Deng et al., 2024; Devanathan and Boyd, 2024). The Polyak stepsize requires knowledge of the optimal function value $f^{\star}$ and is defined as
@@
PolyakGD automatically adjusts the stepsize according to the position of the current iterate $x^{k}$ on the optimization landscape. The resulting stepsizes are typically much larger than standard constant or diminishing stepsizes, making PolyakGD more aggressive and often yielding superior practical performance (Zamani and Glineur, 2024). Consequently, the Polyak stepsize is widely used in settings where the optimal function value $f^{\star}$ is known, such as convex feasibility problems and over-parameterized machine learning models (Boyd, 2013; Loizou et al., 2021). Even when $f^{\star}$ is not known a priori, PolyakGD remains highly effective, provided that a reasonable estimate of the optimal value is available (Polyak, 1969; Hazan and Kakade, 2019; Zheng et al., 2024).
@@
From a theoretical perspective, the convergence of the Polyak stepsize has been extensively studied in nonsmooth convex optimization. Polyak (1987) proved that the subgradient method with the Polyak stepsize satisfies $\lim_{k\to\infty}\sqrt{k}\,(f(x^{k})-f^{\star})=0$. Moreover, the convergence rate of the best-iterate [^1] By best-iterate, we refer to $\min_{1\leq k\leq K}(f(x^{k})-f^{\star})$; by last-iterate, we refer to $f(x^{K})-f^{\star}$. is $\mathcal{O}(1/\sqrt{K})$ (see, e.g., Boyd (2013)). Beyond best-iterate guarantees, Zamani and Glineur (2024) analyzed the last-iterate behavior of the Polyak stepsize, establishing an $\mathcal{O}(1/\sqrt[4]{K})$ convergence rate and providing a worst-case function to match this rate. Their analysis uses the methodology based on performance estimation problem (PEP) (Drori and Teboulle, 2014; Taylor et al., 2017), building on earlier work (Zamani and Glineur, 2025) [^2] Zamani and Glineur (2024, 2025) employ PEP to obtain numerical dual multipliers that guide the derivation of analytical expressions for their proofs.. Additionally, Criscitiello and Boumal (2023) provided a worst-case function for subgradient method with the Polyak stepsize on Riemannian manifolds. More recently, Orabona and D’Orazio (2025) showed that the Polyak stepsize can be interpreted as applying the subgradient method to a specific surrogate function, providing an alternative justification for the necessity of acquiring $f^{\star}$. In contrast to the well-studied nonsmooth setting, the smooth setting remains largely unexplored. The primary reference is the note Hazan and Kakade (2019) by Hanzan and Kakade, who established convergence guarantees for PolyakGD on both smooth convex and smooth strongly convex objectives. In the smooth convex setting, they proved an $\mathcal{O}(1/K)$ convergence rate for the best-iterate, matching the rate of constant-stepsize gradient descent.
Although PolyakGD is adaptive and often exhibits superior practical performance, the existing $\mathcal{O}(1/K)$ convergence rate does not exceed that of gradient descent with a constant stepsize. It naturally motivates the following question:
@@
Two main approaches to tight convergence analysis have been developed in the literature. One is to construct worst-case functions. For example, Drori and Teboulle (2014); Teboulle and Vaisbourd (2023) designed the Huber loss to justify the tightness of the $\mathcal{O}(1/K)$ rate for constant-stepsize gradient descent. The other one is PEP, which works well for designing a fixed stepsize schedule (Altschuler and Parrilo, 2025; Grimmer, 2024; Rubbens et al., 2025). To our best knowledge, there are only two papers on the tight convergence of the Polyak stepsize, motivated from the PEP-based methodology. They focus on the strongly convex case and study the variants of the original Polyak stepsize (Barré et al., 2020; Huang and Qi, 2024). PEP remains applicable in the strongly convex setting, as the contraction ratio between consecutive iterations can be explicitly analyzed. However, when the objective function is not strongly convex, PEP does not readily extend to adaptive stepsizes (Goujaud et al., 2023, 2024).
Our paper provides a two-fold answer to the above question. On the one hand, we show that the known convergence rates for PolyakGD are indeed tight by constructing worst-case functions. On the other hand, we demonstrate that under floating-point arithmetic, PolyakGD can automatically leverage numerical error to escape this worst-case behavior. This interesting phenomenon explains the superior empirical performance of PolyakGD in practice. Theoretically, the strong performance of PolyakGD is attributed to its inherent adaptivity, a crucial property for modern optimization methods (Malitsky and Mishchenko, 2020; Gao et al., 2024; Li and Lan, 2025; He et al., 2025)Hazan and Kakade (2019) showed that the Polyak stepsize is simultaneously adaptive to both smooth and nonsmooth regimes, encompassing general convex as well as strongly convex objectives. Mishkin et al. (2024) further established that it adapts to directional smoothness. More recently, Orabona and D’Orazio (2025) demonstrated that the Polyak stepsize also adapts to the curvature of a surrogate function induced by the original objective. Inspired by these results, we further advance this line of research and provide a comprehensive answer to the following question:
@@
Contributions
This work answers the above two questions. To address the first question, we construct a simple two-dimensional quadratic function. Our proof strategy relies on choosing a specific initial point such that the Polyak stepsize reduces to a constant stepsize along the trajectory in the strongly convex case. Subsequently, we derive the worst-case functions for the general convex and Hölder smooth cases by adapting this quadratic function. Based on this construction, we establish the tightness of the linear convergence rate for strongly convex functions, the $\mathcal{O}(1/K)$ rate for general convex functions, and the $\mathcal{O}(1/K^{(\nu+1)/2})$ rate for $\nu$-Hölder smooth objectives. We then show that under floating-point arithmetic, PolyakGD escapes this worst-case behavior.
We manage to address the second question by demonstrating that the Polyak stepsize is universal across various function classes. We establish that PolyakGD simultaneously adapts to both Hölder smoothness and Hölder growth conditions (Theorem 4.1). In particular, when only the Hölder growth condition holds, PolyakGD achieves the optimal convergence rate; when the function is Hölder smooth, it matches the rate of the universal gradient method of Nesterov (2015). Furthermore, we show that the Polyak stepsize adapts to the recently proposed intrinsic global curvature bound Nesterov (2025), and that our universal analysis extends to the stochastic setting under the interpolation condition. A summary of these results is provided in Table 1.
Table 1: Upper/lower bounds of PolyakGD.
@@
Hazan and Kakade (2019) Theorem 3.2
@@
Hazan and Kakade (2019) Theorem 3.1
@@
Theorem 4.1 Nemirovskii and Nesterov (1985)
@@ 2 Preliminaries
where the objective function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is convex and differentiable [^3] Our analysis can be extended to the projected (sub)gradient method for the constrained setting $\min_{x\in\mathcal{X}}f(x)$ with a closed convex constraint set $\mathcal{X}$ and $f$ is nonsmooth.. We assume that the optimal solution set $\mathcal{X}^{\star}$ is nonempty and the optimal function value $f^{\star}=f(x^{\star})$ is finite. Besides, we assume that $f^{\star}$ is known. Addressing the scenario where $f^{\star}$ is unknown is beyond the scope of this paper, and the techniques developed in Polyak (1969); Hazan and Kakade (2019) can be applied to estimate $f^{\star}$ at the cost of an additional $\log(1/\varepsilon)$ factor in the overall complexity.
@@
with proofs provided in Orabona (2023). For the lower curvature bound, we adopt the following growth condition:
@@
Consequently, several subtle points deserve attention. First, we require that $\mathcal{X}^{\star}\subseteq\mathcal{K}$, otherwise the Hölder growth condition does not seem intuitive. As a result, we exclude the case $r<\nu+1$, as choosing $x$ sufficiently close to $\mathcal{X}^{\star}$ would lead to a contradiction with the inequality (6). Moreover, we require that $\sup_{x\in\mathcal{K}}\operatorname{dist}(x,\mathcal{X}^{\star})<\infty$. Otherwise, the case $r>\nu+1$ would also yield a contradiction by taking $x$ such that $\operatorname{dist}(x,\mathcal{X}^{\star})\rightarrow\infty$ (see Section 1.2 of Roulet and d’Aspremont (2017) for a similar argument). Fortunately, as discussed at the beginning of Section 4, $\gamma$-PolyakGD (Algorithm 1) naturally ensures the existence of such a proper set $\mathcal{K}$ due to the Fejér monotonicity.
@@ 3 Tight Convergence of $\gamma$-PolyakGD
Consequently, the main task is to construct a worst-case function to answer the above question. For instance, Drori and Teboulle (2014); Teboulle and Vaisbourd (2023) utilized the Huber loss to establish the tightness of the $\mathcal{O}(1/K)$ rate for constant stepsize gradient descent. The intuition is that the small gradient magnitude of the Huber loss forces a constant stepsize gradient descent to take conservatively small steps, thereby justifying the tightness. However, this strategy fails due to the adaptivity of $\gamma$-PolyakGD, as demonstrated below.
@@ 3.1 Worst-case quadratic functions for $\gamma$-PolyakGD
Our reduction argument is also applicable to the BB stepsize on quadratic functions due to its scale invariance. By leveraging the tight convergence result and the worst-case quadratic function from Li and Sun (2021), a similar $\Omega(1/K)$ result can be established for the BB stepsize.
@@
Setting $\nu\rightarrow 0$, the above result reduces to the lower bound $\Omega(1/\sqrt{K})$. We remark that this result does not contradict the $\Omega(1/\sqrt[4]{K})$ bound established in Zamani and Glineur (2024). Our tightness result is established for the best iterate, while theirs applies to the last iterate, which is generally a more challenging criterion. Consequently, the worst-case function constructed in Zamani and Glineur (2024) is more complicated. This distinction reveals the significant gap between the best iterate and the last iterate under bounded gradient condition, highlighting the necessity of analyzing both cases.
@@
Establishing the tight gradient norm convergence of constant stepsize gradient descent involves subtle distinctions. Setting $\nu=1$, the above result reduces to the lower bound $\Omega(1/\sqrt{K})$. However, Nesterov and others (2018); Diakonikolas and Wang (2021); Teboulle and Vaisbourd (2023) established that small constant stepsizes achieve an upper bound of $\mathcal{O}(1/K)$, a result shown to be tight by Kim and Fessler (2018) via PEP. Using a slightly large constant stepsize slows the convergence to $\mathcal{O}(1/\sqrt{K})$, as shown in Proposition 4.4 in Rotaru et al. (2022) and Theorem 2.1 in Rotaru et al. (2024), a bound that is also tight according to PEP. Therefore, our result suggests that the Polyak stepsize behaves like a “large constant stepsize” in some sense.
@@ 3.2 Polyak stepsize escapes the worst case under floating-point arithmetic
Our tightness convergence results show that PolyakGD reduces to gradient descent with a constant stepsize on the worst-case quadratic function. While this conclusion holds under exact arithmetic, it may not persist in practical implementations due to floating-point errors. Indeed, when implementing the Polyak stepsize on the constructed worst-case quadratic function, the numerical behavior differs from the theoretical results. Figure 2 plots the stepsize sequence $\{\alpha_{k}\}$ and the corresponding function value gaps $\{q(x^{k})-q^{\star}\}$ generated by $1$-PolyakGD. As predicted by our theory, $1$-PolyakGD initially behaves like constant stepsize gradient descent. Yet after a few iterations, it departs from the theoretical trajectory, producing larger stepsizes and converging significantly faster. This suggests that under inexact arithmetic, $\gamma$-PolyakGD (with $\gamma\in(0,2)$) can automatically exploit floating-point errors to escape the worst-case behavior, which may explain the superior practical performance of the Polyak stepsize. This section formally quantifies this intuition. The techniques developed in this section have also been applied to analyze the hypergradient descent heuristic recently (Chu et al., 2025).
@@
and $s_{1}$ (or equivalently $s_{2}$) is a period-2 orbit of $\mathcal{F}_{\gamma}$ (Alligood et al., 1997). To understand the local behavior of $\gamma$-PolyakGD in the presence of floating-point errors, it is natural to examine the stability of this nonlinear map around its orbit. This reduces to analyzing the spectrum of the Jacobian of $\mathcal{F}_{\gamma}$. The proof of Proposition 3.2 is assisted by Mathematica and the source code is available at this link.
@@
Proposition 3.2 shows that when $\gamma=1$ and $\kappa>6$, the spectral radius of the product Jacobian is strictly greater than $1$. Hence, the nonlinear dynamical system (13) is unstable around the trajectory induced by the worst-case function. Consequently, once floating-point errors enter the iterations, $\gamma$-PolyakGD deviates from this trajectory and exhibits accelerated convergence. Similar observation holds for all $\gamma\in(0,2)$. In contrast, when $\gamma=2$, the spectral radius is exactly $1$, implying that $2$-PolyakGD may become trapped on the worst-case trajectory. It is consistent with the zig-zag behavior observed in Huang and Qi (2024). In summary, the worst-case trajectory is only realized under exact arithmetic, and PolyakGD in practice is largely immune to it.